OpenStudy (anonymous):

If the ratio of the volumes of two similar solids is 8:125, then what is the ratio of their surface areas?

OpenStudy (amistre64):

well, we can compare the volumes of a cube ot that of a square ...

OpenStudy (amistre64):

a cube of side cbrt(8) and a cube of side sqrt(125) the squares have areas of (cbrt(8))^2 and (cbrt(125))^2

OpenStudy (amistre64):

what about cones? V = 4/3 pi r^3 8V = 4/3 pi (r cbrt(8))^3 A = 4 pi (r cbrt(8))^2 same ratio

OpenStudy (anonymous):

I still don't understand.

OpenStudy (amistre64):

the squares have areas of (cbrt(8))^2 and (cbrt(125))^2 what more is there to say?

OpenStudy (anonymous):

I just don't get it. Why are you using squares?

OpenStudy (anonymous):

@jim_thompson5910 You're good at explaining things.. Do you know the answer?